Calculate the mole fraction and molality of glycer Please explain
Answer #1

mole fractin = mol of solute / total mol

mol of solute = mass/MW = 45/92.09 = 0.4886 mol

mol of solvent = mas/MW = 100/18 = 5.55 mol

total mol = 0.4886 + 5.55 = 6.0386 mol

frac = 0.4886/6.0386 = 0.0809127

molality

m = mol solut e/ kg solvent = 0.4886 /(0.1) = 4.886 m